What is the slope of the line 3/4x+1/4y=1/2
1
Expert's answer
2012-10-23T08:32:10-0400
Equation of a straight line in slope-intercept form looks like following
y=kx+b where k-slope
So let's transform given equation into slope-intercept form:
3/4x+1/4y=1/2
3x+y=2
y=-3x+2
So slope k=-3

For further information you can watch our video about equation of a line: